# Two-dimensional inference of divertor plasma characteristics: advancements to a multi-instrument Bayesian analysis system

### v1

@@ 5.2.1 Importance of including PMI in the Balmer emission model
Posterior ii only considers atomic contributions to Balmer line emission which results in a sustained overestimation of the electron temperature outside of the region where electron-ion recombination emission is dominant as can be seen in figure 5(a). The cause of this was that the Balmer line emission is increased by the presence of plasma-molecule interaction. Because the Balmer line ratios from PMI and EIE are similar and significantly different from EIR, these brighter images could not be explained by EIR. Therefore, the model omitting PMI was compelled to attribute these bright images to EIE which, in turn, demanded higher electron temperatures, leading to the overestimate. This overestimation is an important demonstration that an incorrect model in Bayesian inference does not necessitate the true solution to fall within the found uncertainty interval if an invalid emission model is applied. This highlights the importance of using a model that accommodates PMI when required.
@@ 5.2.2 Importance of Spatially-Dependent Priors
The spatially-dependent priors are permitted through jointly conducting inference with parameters at each mesh vertex as allowed by the mesh-based inference paradigm. Figure 5(b) demonstrates that the omission of these in posterior iii, substantially increased uncertainty on the electron density (which was also observed for other inferred parameters) and the complete loss of information out of view of the synthetic camera diagnostic (as expected). The lack of additional information on trends in static electron pressure and smoothness leaves an ill-posed problem in which many different combinations of parameters can accurately recreate the multi-wavelength imaging data thus causing large uncertainty in the inference. This highlights the benefit of jointly considering many parameters at different spatial positions in the inference as permitted by the mesh-based D-MIBAS.
@@ 5.2.3 Importance of helium emission
Posterior iv reflects situations where helium data are unavailable or unreliable. The helium singlet lines are known to carry information on both the electron temperature and electron density. The exclusion of helium data degraded the MdAPE for $T_{\mathrm{e}}$ and $n_{\mathrm{e}}$ to 14% and 11% respectively as shown in Figure 5.
Much of the reduction in the error of parameters inferred by posterior i over iv was observed at low electron temperatures and densities away from the separatrix. Along the separatrix, the inferred MAP estimate of posterior iv was in good agreement with the true plasma values; however, the inclusion of helium data in posterior i reduced the uncertainty in the inference of $T_{\mathrm{e}}$ and $n_{\mathrm{e}}$. As shown in figure 5(c), a consequence of this was that the inclusion of helium data was found to significantly improve the $n_{\mathrm{H}}$ inference along the separatrix despite the $n_{\mathrm{H}}$ parameter not appearing in the helium forward models. This was because the reduced uncertainty in the inference of $T_{\mathrm{e}}$ and $n_{\mathrm{e}}$ managed to better constrain the source of the Balmer-line emission (EIR/PMI/EIE). By having a good understanding of how much EIE emission (which is $n_{\mathrm{H}}$ dependent) occurs, the $n_{\mathrm{H}}$ inference improved. This serves to demonstrate how the inclusion of additional diagnostics in the IDA can improve the accuracy of the inference of certain parameters even if they themselves are not involved in the forward model of the additional diagnostic.
@@
Figure 6: Comparison of inferred parameters with the ‘true’ values (figure 2Ia) for the detached case. The parameters follow a surface of constant poloidal flux, $\psi_{N}=1$ (the separatrix).
@@ 6 Discussion
First, we will discuss the limitations of the current analysis and possible improvements for the future with an emphasis on the preparation of applying D-MIBAS to experimental data. Afterwards, we will discuss possible applications of this analysis and what the implications are for such techniques.
@@ 6.1.1 Consideration for experimental data application: forward model inaccuracies
The diagnostic description of imaging and spectroscopic data include multiple uncertainties. Multi-wavelength imaging employs bandpass filters to monitor a small, spectrally integrated, part of the emission spectra. Plasma background emission and neighbouring spectral transitions can contaminate the monitored signal, resulting in overestimated camera brightnesses [47]. Reflections can displace the monitored emission which, if present, would require a more complex diagnostic description [26]. Similarly, photon opacity [48], if present, would require a substantial modification to the diagnostic description. Additionally, there are uncertainties in the spatial and absolute calibrations of imaging and spectroscopy systems. The various photon emission coefficients used in the emission models can also have significant uncertainties of 10-20 % [11].
The additional uncertainties outlined above requires that, when applied to experimental data, the forward model uncertainties (held in $\Sigma$ of (3)) will exceed those from the diagnostic measurements alone. In turn, the uncertainties on inferred parameters will increase from those observed in Figure 6. To limit the impact of the increase in forward model uncertainties, there are several methods that can be used:
@@
  • Monte Carlo approaches can be used to perturb uncertain quantities within reasonable intervals (e.g., perturb spatial camera calibration uncertainties [21]) to create multiple posterior distributions that can be combined.
@@
D-MIBAS provides a framework for comparing the consistency between the data obtained by the different diagnostics. This makes integrated data analysis particularly suitable for investigating errors in diagnostic descriptions (forward models). This can be strengthened by including additional diagnostics in the D-MIBAS inference (section 6.1.2). Awareness of inadequate diagnostic descriptions can be used to improve instrument characterisation [20, 21] (e.g., improved calibrations) and to inform where more comprehensive models are required (e.g., modelling directly to the camera image data instead of pre-inverted data [23]).
@@ 6.1.2 Improving D-MIBAS analysis through additional diagnostics
The inclusion of additional diagnostics has multiple benefits. They can: reduce parameter inference uncertainty and improve inference accuracy; mitigate model uncertainties; expose deficiencies in diagnostic models; and act to validate diagnostic interpretation. However, care must be taken that the inclusion of an additional diagnostic does not result in too large of an increase in the number of free parameters required. Examples of additional diagnostics that can be included in the current D-MIBAS setup without requiring additional free parameters include:
@@
  • Bolometry (including higher resolution imaging bolometry [49]) can provide additional information on the total radiation. This provides an upper-limit to the permitted hydrogenic radiation (of which can be modelled from the current parameter set).
@@ 6.2 Implication and relevance of D-MIBAS for furthering exhaust understanding
The divertor region complexities requires the use of information from a wide range of diagnostics. The implementation of D-MIBAS in this work makes advances in all four of these areas:
@@ 6.2.1 Derived Quantities
Further derived parameters can be obtained from electron and neutral parameter information. The multi-dimensional posterior distribution provides correlated samples which can be used to compute derived parameters, including their uncertainty. The inferred parameter set can be used to obtain: static electron pressure [12]; heat and ion fluxes parallel to flux tubes [11]; estimates of ion sources and sinks (both due to atomic (EIE, EIR) as well as molecular processes (MAR, MAI)) [44] and (hydrogenic) power loss estimates [50, 44]. This technique was utilised to compute the EIR, EIE and PMI emission fractions for $D\gamma$ in figure 8. In figure  7, the derived parallel heat flux and static electron pressure is shown at a roughly constant magnetic connection length to the target in the vicinity of the detached case’s detachment front. Their associated uncertainties are considerable, however, due to the compounding of uncertainties from multiple parameters.
@@ 6.3 Relevance of D-MIBAS for reactors
Reactors will have to operate with a reduced set of diagnostics due to the harsher reactor conditions (e.g. a neutron-rich environment, intense heat fluxes, limitation to robotic access) and port space will be limited [1]. Reactor diagnostics may fail and become redundant during operation. D-MIBAS provides a robust statistical framework for investigating how much information a set of diagnostics can provide. This can be used to find the diagnostic coverage required to produce a reactor’s acceptable level of information on the divertor plasma state and the impact of losing a diagnostic from that set. Langmuir probes measure ion fluxes to chamber walls which is widely used in the community to quickly interpret the plasma state and to understand the power being received by the chamber walls. However, Langmuir probes are unlikely to be operational in harsh reactor conditions. So long as inferences can still be made with sufficient accuracy, the D-MIBAS framework can still provide such quantities even in the absence of the diagnostic.
In reactors, real-time control of the power exhaust problem is required. This requires understanding what information from the divertor (for example, detachment front position) will need to be controlled in real-time. The insight provided by applying D-MIBAS to experimental data can be used to answer that question. Although D-MIBAS is too numerically intensive to be performed in real-time currently, we envisage it to be an important aid (in training, validation and uncertainty quantification) for neural-network approaches to predicting the divertor state.
Reactor conditions will be significantly different from the MAST-U Super-X plasma conditions and synthetic diagnostics used to test the current D-MIBAS methodology. This will likely require modifications to our forward models such as the consideration of photon opacity in our emission models [48]. Although such D-MIBAS analysis tools hold a strong potential for reactors, substantial effort will be required to test such approaches in reactor-relevant divertor conditions using a reactor-relevant diagnostic set. However, as outlined in this work, the insight into D-MIBAS’s abilities and limitations can be gleaned through synthetic data without needing to wait for reactors to begin operation.
@@ A.1 Molecular Emission Model
Most Balmer line emission arising from molecular break-up in the divertor occurs when molecular ions (mostly $D_{2}^{+}$, but also $D_{2}^{-}\rightarrow D^{-}+D$ and $D_{3}^{+}$) interact with the plasma [51]. Modeling the Balmer line emission from plasma-molecular interactions requires the photon-emissivity coefficients  [^6] Akin to those provided by ADAS for atomic processes as used in. for each process, $\mathrm{PEC}{}_{D_{2}}$, $\mathrm{PEC}{}^{n\to 2}_{D_{2}^{+}}$, $\mathrm{PEC}{}^{n\to 2}_{D^{-}}$, $\mathrm{PEC}{}^{n\to 2}_{D_{3}^{+}}$, as well as each of the respective molecular densities.
@@
This encapsulates the molecular contribution to the Balmer-line emissivity (with each $\mathrm{PEC}{}$ term depending on $n_{\mathrm{e}}$ and $T_{\mathrm{e}}$ alone) whilst avoiding the need to include additional parameters for each molecular process in  (5).
@@
Under the assumption of no transport, the ratios of the various molecular ion densities to the molecular density 777assuming the electron density and the hydrogen ion density are roughly equal in (10) can be estimated by the ratio between the creation and destruction rates of each molecular species [44, 11]. These creation and destruction rates were adopted from the default hydrogenic rates used by Eirene [52] as a function of $T_{\mathrm{e}}$ and $n_{\mathrm{e}}$ [52]. These rates are for hydrogen, but we aim to analyse data of a deuterium plasma. As such, translation factors for hydrogen to deuterium, adopted from [53], were applied ($0.95$ and $0.7$ for $\frac{n_{D_{2}^{+}}}{n_{D_{2}}}$ and $\frac{n_{D^{-}}}{n_{D_{2}}}$ respectively), rather than using an isotope mass-based rescaling, which exacerbates errors in the molecular charge-exchange rates at low temperatures [46]. The respective $\mathrm{PEC}{}$ values for plasma-molecular interactions were obtained from the population coefficients retrieved from the YACORA-on-the-web [54, 55] collisional-radiative model. In detachment-relevant conditions (0.6-2 eV), $D_{2}^{+}$ is the dominant driver of hydrogen emission associated with plasma-molecular break-up: $\mathrm{PEC}{}_{D{2,\mathrm{eff}}{}}^{n\rightarrow 2}(T_{e},n_{e})\approx\frac{n_{D_{2}^{+}}}{n_{D_{2}}}\mathrm{PEC}{}_{D_{2}^{+}}^{n\rightarrow 2}$.
@@
The uncertainties in the relative ratios of molecular ion densities to $D_{2}$ in (10) are large [8]. There are strong concerns about the various creation rates for molecular ions that Eirene employs, particularly $D_{2}+D^{+}\rightarrow D_{2}^{+}+D$ which is the dominant driver for $D_{2}^{+}$ in detached conditions [11, 28, 46, 50, 27, 56]. Molecular ion creation is driven by vibrationally excited molecules, which depend not only on local $n_{\mathrm{e}}$ and $T_{\mathrm{e}}$, but also on molecular transport and plasma-wall interactions. Additionally, there are strong concerns about the vibrational distribution model employed by Eirene [11, 46, 27]. These concerns lead to uncertainties of multiple orders of magnitude on the amplitude of $\mathrm{PEC}{}_{D{2,\mathrm{eff}}{}}^{n\rightarrow 2}$.
@@
If only one emission process is dominant, the ratio between the emission coefficient of two Balmer line is not affected by these uncertainties  [^8] Due to the uncertainties associated with the molecular ion fractions, we have also performed IDA under the assumption that $D_{2}^{+}$ is the dominant emission process: $\frac{\mathrm{PEC}{}_{D{2,\mathrm{eff}}{}}^{n_{1}\rightarrow 2}}{\mathrm{PEC}{}_{D{2,\mathrm{eff}}{}}^{n_{2}\rightarrow 2}}\approx\frac{\mathrm{PEC}{}_{D_{2}^{+}}^{n_{1}\rightarrow 2}}{\mathrm{PEC}{}_{D_{2}^{+}}^{n_{2}\rightarrow 2}}$, which did not change the outcome of our analysis within uncertainties.. Therefore, rather than multiplying $\mathrm{PEC}{}_{D{2,\mathrm{eff}}{}}^{n\to 2}$ by $n_{\mathrm{e}}$ and an effective molecular density, we express our emission model in such a way that only the ratios of $\mathrm{PEC}{}_{D{2,\mathrm{eff}}{}}^{n\rightarrow 2}$ between different Balmer lines are used. This re-parameterisation utilises $Q_{\mathrm{mol}}$ which is defined to be the ratio of $D_{\alpha}$’s molecular emissivity contributions, $\varepsilon^{3\to 2}_{\mathrm{mol}}$, to atomic emissivity contributions, $\varepsilon^{3\to 2}_{\mathrm{atm}}$,
@@
$\eqalign{\varepsilon_{n\to 2}=\underbrace{n_{e}^{2}\mathrm{PEC}{}^{n\to 2}_{rec}+n_{0}n_{e}\mathrm{PEC}{}^{n\to 2}_{\mathrm{exc}}}_{\mathrm{Atomic\;contribution,}\;\varepsilon^{n\to 2}_{\mathrm{atm}}}+\cr\ \underbrace{Q_{D_{\alpha}}^{\mathrm{mol}}\left(n_{e}^{2}\mathrm{PEC}{}^{3\to 2}_{\mathrm{rec}}+n_{0}n_{e}\mathrm{PEC}{}^{3\to 2}_{\mathrm{exc}}\right)\frac{\mathrm{PEC}{}_{D{2,\mathrm{eff}}{}}^{n\to 2}}{\mathrm{PEC}{}_{D{2,\mathrm{eff}}{}}^{3\to 2}}}_{\mathrm{Molecular\;contribution,}\;\varepsilon^{n\to 2}_{\mathrm{mol}}}.}$ (12)

### v2

@@ 5.2.1 Importance of including PMI in the Balmer emission model
Posterior ii only considers atomic contributions to Balmer line emission which results in a sustained overestimation of the electron temperature outside of the region where electron-ion recombination emission is dominant as can be seen in Figure 6(a). The cause of this was that the Balmer line emission is increased by the presence of plasma-molecule interaction. Because the Balmer line ratios from PMI and EIE are similar and significantly different from EIR, these brighter images could not be explained by EIR. Therefore, the model omitting PMI was compelled to attribute these bright images to EIE which, in turn, demanded higher electron temperatures, leading to the overestimate. This overestimation is an important demonstration that an incorrect model in Bayesian inference does not necessitate the true solution to fall within the found uncertainty interval if an invalid emission model is applied. This highlights the importance of using a model that accommodates PMI when inferring divertor plasmas with substantial molecular hydrogen densities.
@@ 5.2.2 Importance of Spatially-Dependent Priors
The spatially-dependent priors are permitted through jointly conducting inference with parameters at each mesh vertex as allowed by the mesh-based inference paradigm. Figure 6(b) demonstrates that the omission of these in posterior iii substantially increased uncertainty on the electron density (which was also observed for other inferred parameters) and the complete loss of information out of view of the synthetic camera diagnostic (as expected). The lack of additional information on trends in static electron pressure and smoothness leaves an ill-posed problem in which many different combinations of parameters can accurately recreate the multi-wavelength imaging data, thus causing large uncertainty in the inference. This highlights the benefit of jointly considering many parameters at different spatial positions in the inference as permitted by the mesh-based D-MIBAS.
@@ 5.2.3 Importance of helium emission
Posterior iv reflects situations where helium data are unavailable or unreliable. The helium singlet lines are known to carry information on both the electron temperature and electron density. The exclusion of helium data degraded the MdAPE for $T_{\mathrm{e}}$ and $n_{\mathrm{e}}$ to 14% and 11% respectively, as shown in Figure 5.
Along the separatrix, the inferred MAP estimate of posterior iv was in good agreement with the true plasma values; however, the inclusion of helium data in posterior i reduced the uncertainty in the inference of $T_{\mathrm{e}}$ and $n_{\mathrm{e}}$ (Figure 5). As shown in Figure 6(c), the inclusion of helium data was found to significantly improve the $n_{\mathrm{H}}$ inference along the separatrix despite the $n_{\mathrm{H}}$ parameter not appearing in the helium forward models. This was because the reduced uncertainty in the inference of $T_{\mathrm{e}}$ and $n_{\mathrm{e}}$ managed to better constrain the process leading to Balmer-line emission (EIR/PMI/EIE). By having a good understanding of how much EIE emission (which is $n_{\mathrm{H}}$ dependent) occurs, the $n_{\mathrm{H}}$ inference improved. This serves to demonstrate how the inclusion of additional diagnostics in the IDA can improve the accuracy of the inference of certain parameters even if they themselves are not involved in the forward model of the additional diagnostic.
@@
Figure 6: Comparison of inferred parameters with the ‘true’ values (Figure 2Ia) for the detached case. The parameters follow a surface of constant poloidal flux, $\psi_{N}=1$ (the separatrix). Solid lines indicate the inferred (MAP) parameters and the shaded region indicate the 95% HDI.
@@ 6 Discussion
In this section, we will discuss the limitations of the current analysis and possible improvements for the future with an emphasis on the preparation of applying D-MIBAS to experimental data. We will go on to discuss possible applications of this analysis and what the implications are for such techniques.
@@ 6.1.1 Consideration for experimental data application: forward model inaccuracies
The diagnostic description of imaging and spectroscopic data includes multiple uncertainties. Multi-wavelength imaging employs bandpass filters to monitor a small, spectrally integrated part of the emission spectra. Plasma background emission and neighbouring spectral transitions can contaminate the monitored signal, resulting in overestimated camera brightnesses [Perek2021]. Reflections can displace the monitored emission which, if present, would require a more complex diagnostic description [Karhunen2023]. Similarly, photon opacity [Pshenov2023, Lomanowski2020, Lawson_2024], if present, would require a substantial modification to the diagnostic description. Additionally, there are uncertainties in the spatial and absolute calibrations of imaging and spectroscopy systems. The various photon emission coefficients used in the emission models can also have significant uncertainties of 10-20% [Verhaegh2021b].
The additional uncertainties outlined above require that, when applied to experimental data, the forward model uncertainties (held in $\Sigma$ of equation (3)) will exceed those from the diagnostic measurements alone. In turn, the uncertainties on inferred parameters will increase from those observed in Figure 6. To limit the impact of the increase in forward model uncertainties, there are several methods that can be used:
@@
  • Monte Carlo approaches can be used to perturb uncertain quantities within reasonable intervals (e.g., perturb spatial camera calibration uncertainties [Perek_2022]) to create multiple posterior distributions that can be combined.
@@
D-MIBAS provides a framework for comparing the consistency between the data obtained by the different diagnostics. This makes integrated data analysis particularly suitable for investigating errors in diagnostic descriptions (forward models). This can be strengthened by including additional diagnostics in the D-MIBAS inference (Section 6.1.2). Awareness of inadequate diagnostic descriptions can be used to improve instrument characterisation [Perek2019submitted, Perek_2022] (e.g., improved calibrations) and to inform where more comprehensive models are required (e.g., modelling directly to the camera image data instead of pre-inverted data [Bowman_2020]).
@@ 6.1.2 Improving D-MIBAS analysis through additional diagnostics
The inclusion of additional diagnostics has multiple benefits. They can: reduce parameter inference uncertainty and improve inference accuracy; mitigate model uncertainties; expose deficiencies in diagnostic models; and act to validate diagnostic interpretation. However, care must be taken that the inclusion of an additional diagnostic does not result in too large an increase in the number of free parameters required. Examples of additional diagnostics that can be included in the current D-MIBAS setup without requiring additional free parameters include:
@@
  • Bolometry (including higher resolution imaging bolometry [Federici2023]) can provide additional information on the total radiation. This provides an upper-limit to the permitted hydrogenic radiation (of which can be modelled from the current parameter set).
@@ 6.2 Implication and relevance of D-MIBAS for furthering exhaust understanding
The divertor region complexities require the use of information from a wide range of diagnostics. The implementation of D-MIBAS in this work makes advances in all four of these areas:
@@ 6.2.1 Derived Quantities
Further derived parameters can be obtained from electron and neutral parameter information. The multi-dimensional posterior distribution provides correlated samples which can be used to compute derived parameters, including their uncertainty. The inferred parameter set can be used to obtain: static electron pressure [Lipschultz2007a]; heat and ion fluxes parallel to flux tubes [Verhaegh2021b]; estimates of ion sources and sinks (both due to atomic (EIE, EIR) as well as molecular processes (MAR, MAI)) [Verhaegh2021a] and (hydrogenic) power loss estimates [Verhaegh2023b, Verhaegh2021a]. This technique was utilised to compute the EIR, EIE and PMI emission fractions for $D\gamma$ in Figure 7. In Figure 8, the derived parallel heat flux and static electron pressure are shown at a roughly constant magnetic connection length to the target in the vicinity of the detached case’s detachment front. Their associated uncertainties are considerable, however, due to the compounding of uncertainties from multiple parameters.
Figure 7: Comparison of true (a) and inferred (b) (using posterior i) dominant emission regions for the Dγ Balmer line for the detached case.
@@ 6.3 Relevance of D-MIBAS for reactors
Reactors will have to operate with a reduced set of diagnostics due to the harsher reactor conditions (e.g. a neutron-rich environment, intense heat fluxes, limitation to robotic access) and port space will be limited [Wenninger2014]. Reactor diagnostics may fail and become redundant during operation. D-MIBAS provides a robust statistical framework for investigating how much information a set of diagnostics can provide. This can be used to find the diagnostic coverage required to produce a reactor’s acceptable level of information on the divertor plasma state and the impact of losing a diagnostic from that set. Langmuir probes measure ion fluxes to chamber walls which are widely used in the community to quickly interpret the plasma state and to understand the power being received by the chamber walls. However, Langmuir probes are unlikely to be operational in harsh reactor conditions. So long as inferences can still be made with sufficient accuracy, the D-MIBAS framework can still provide such quantities even in the absence of the diagnostic.
In reactors, real-time control of the power exhaust problem is required. This requires understanding what information from the divertor (for example, detachment front position) will need to be controlled in real-time. The insight provided by applying D-MIBAS to experimental data can be used to answer that question. Although D-MIBAS is too numerically intensive to be performed in real-time currently, we envisage it to be an important aid (in training, validation, and uncertainty quantification) for neural-network approaches to predicting the divertor state.
Reactor conditions will be significantly different from the MAST-U Super-X plasma conditions and synthetic diagnostics used to test the current D-MIBAS methodology. This will likely require modifications to our forward models, such as the consideration of photon opacity in our emission models [Pshenov2023]. Although such D-MIBAS analysis tools hold a strong potential for reactors, substantial effort will be required to test such approaches in reactor-relevant divertor conditions using a reactor-relevant diagnostic set. However, as outlined in this work, the insight into D-MIBAS’s abilities and limitations can be gleaned through synthetic data without needing to wait for reactors to begin operation.
@@ A.1 Molecular Emission Model
Most Balmer line emission arising from molecular break-up in the divertor occurs when molecular ions (mostly $D_{2}^{+}$, but also $D_{2}^{-}\rightarrow D^{-}+D$ and $D_{3}^{+}$) interact with the plasma [atoms4040026]. Modeling the Balmer line emission from plasma-molecule interactions requires the photon-emissivity coefficients  [^7] Akin to those provided by ADAS for atomic processes as used in equation5. for each process, $\mathrm{PEC}{}_{D_{2}}$, $\mathrm{PEC}{}^{n\to 2}_{D_{2}^{+}}$, $\mathrm{PEC}{}^{n\to 2}_{D^{-}}$, $\mathrm{PEC}{}^{n\to 2}_{D_{3}^{+}}$, as well as each of the respective molecular densities.
@@
This encapsulates the molecular contribution to the Balmer-line emissivity (with each $\mathrm{PEC}{}$ term depending on $n_{\mathrm{e}}$ and $T_{\mathrm{e}}$ alone) whilst avoiding the need to include additional parameters for each molecular process in equation (5).
@@
Under the assumption of no transport, the ratios of the various molecular ion densities to the molecular density 888assuming the electron density and the hydrogen ion density are roughly equal in equation (10) can be estimated by the ratio between the creation and destruction rates of each molecular species [Verhaegh2021a, Verhaegh2021b]. These creation and destruction rates were adopted from the default hydrogenic rates used by Eirene [amjuel] as a function of $T_{\mathrm{e}}$ and $n_{\mathrm{e}}$ [amjuel]. These rates are for hydrogen, but we aim to analyse data of a deuterium plasma. As such, translation factors for hydrogen to deuterium, adopted from [Reiter2018], were applied ($0.95$ and $0.7$ for $\frac{n_{D_{2}^{+}}}{n_{D_{2}}}$ and $\frac{n_{D^{-}}}{n_{D_{2}}}$ respectively), rather than using an isotope mass-based rescaling, which exacerbates errors in the molecular charge-exchange rates at low temperatures [Verhaegh2023a]. The respective $\mathrm{PEC}{}$ values for plasma-molecular interactions were obtained from the population coefficients retrieved from the YACORA-on-the-web [yacora, Wuenderlich2016] collisional-radiative model. In detachment-relevant conditions (0.6-2 eV), $D_{2}^{+}$ is the dominant driver of hydrogen emission associated with plasma-molecular break-up: $\mathrm{PEC}{}_{D{2,\mathrm{eff}}{}}^{n\rightarrow 2}(T_{e},n_{e})\approx\frac{n_{D_{2}^{+}}}{n_{D_{2}}}\mathrm{PEC}{}_{D_{2}^{+}}^{n\rightarrow 2}$.
@@
The uncertainties in the relative ratios of molecular ion densities to $D_{2}$ in equation (10) are large [Verhaegh_2023]. There are strong concerns about the various creation rates for molecular ions that Eirene employs, particularly $D_{2}+D^{+}\rightarrow D_{2}^{+}+D$ which is the dominant driver for $D_{2}^{+}$ in detached conditions [Verhaegh2021b, Verhaegh_2022, Verhaegh2023a, Verhaegh2023b, Verhaegh2023c, Kukushkin2017]. Molecular ion creation is driven by vibrationally excited molecules, which depend not only on local $n_{\mathrm{e}}$ and $T_{\mathrm{e}}$, but also on molecular transport and plasma-wall interactions. Additionally, there are strong concerns about the vibrational distribution model employed by Eirene [Verhaegh2021b, Verhaegh2023a, Verhaegh2023c]. These concerns lead to uncertainties of multiple orders of magnitude in the amplitude of $\mathrm{PEC}{}_{D{2,\mathrm{eff}}{}}^{n\rightarrow 2}$.
@@
If only one emission process is dominant, the ratio between the emission coefficient of two Balmer lines is not affected by these uncertainties . Therefore, rather than multiplying $\mathrm{PEC}{}_{D{2,\mathrm{eff}}{}}^{n\to 2}$ by $n_{\mathrm{e}}$ and the neutral molecular density $n_{D_{2}}$, we express our emission model in such a way that only the ratios of $\mathrm{PEC}{}_{D{2,\mathrm{eff}}{}}^{n\rightarrow 2}$ between different Balmer lines are used. This re-parameterisation utilises $Q_{\mathrm{mol}}$ which is defined to be the ratio of $D_{\alpha}$’s molecular emissivity contributions, $\varepsilon^{3\to 2}_{\mathrm{mol}}$, to atomic emissivity contributions, $\varepsilon^{3\to 2}_{\mathrm{atm}}$,
@@
$\eqalign{\varepsilon_{n\to 2}=\underbrace{n_{e}^{2}\mathrm{PEC}{}^{n\to 2}_{rec}+n_{H}n_{e}\mathrm{PEC}{}^{n\to 2}_{\mathrm{exc}}}_{\mathrm{Atomic\;contribution,}\;\varepsilon^{n\to 2}_{\mathrm{atm}}}+\cr\ \underbrace{Q_{D_{\alpha}}^{\mathrm{mol}}\left(n_{e}^{2}\mathrm{PEC}{}^{3\to 2}_{\mathrm{rec}}+n_{H}n_{e}\mathrm{PEC}{}^{3\to 2}_{\mathrm{exc}}\right)\frac{\mathrm{PEC}{}_{D{2,\mathrm{eff}}{}}^{n\to 2}}{\mathrm{PEC}{}_{D{2,\mathrm{eff}}{}}^{3\to 2}}}_{\mathrm{Molecular\;contribution,}\;\varepsilon^{n\to 2}_{\mathrm{mol}}}.}$ (12)